OpenStudy (anonymous):

does anyone know what the square root of 1/121 is

OpenStudy (mathstudent55):

1^2 = 1 2^2 = 4 3^2 = 9 ... 9^2 = 81 10^2 = 100 11^2 = 121 12^2 = 144 Do you notice that 121 is a perfect square?

OpenStudy (anonymous):

nope not really

OpenStudy (mathstudent55):

1^2 = 1 <--------- 1 is 1^2 2^2 = 4 3^2 = 9 ... 9^2 = 81 10^2 = 100 11^2 = 121 <----------- 121 is 11^2 12^2 = 144 Do you notice that 121 is a perfect square?

OpenStudy (mathstudent55):